OpenStudy (anonymous):

describe how to transform the graph of f into the graph of g. f(x)= sqx and g(x) = sq -x . options are . 1) reflect the graph of f across the y axis. 2)the two graphs are the same. 3)Reflect the graph of f across the y-axis and then reflect across the x-axis. 4)reflect the graph of f across the x axis.

OpenStudy (unklerhaukus):

\[f(x)=x^2\qquad\qquad g(x)=-x^2\] have i got this right?

OpenStudy (anonymous):

i have to choose from the four options.

OpenStudy (unklerhaukus):

do you know what \(y=x^2\) looks like on a x-y chart?